### 08 Feb 2024 v1

# Mamba-ND: Selective State Space Modeling for Multi-Dimensional Data

Abstract
In recent years, Transformers have become the de-facto architecture for sequence modeling on text and a variety of multi-dimensional data, such as images and video. However, the use of self-attention layers in a Transformer incurs prohibitive compute and memory complexity that scales quadratically w.r.t. the sequence length. A recent architecture, Mamba, based on state space models has been shown to achieve comparable performance for modeling text sequences, while scaling linearly with the sequence length. In this work, we present Mamba-ND, a generalized design extending the Mamba architecture to arbitrary multi-dimensional data. Our design alternatively unravels the input data across different dimensions following row-major orderings. We provide a systematic comparison of Mamba-ND with several other alternatives, based on prior multi-dimensional extensions such as Bi-directional LSTMs and S4ND. Empirically, we show that Mamba-ND demonstrates performance competitive with the state-of-the-art on a variety of multi-dimensional benchmarks, including ImageNet-1K classification, HMDB-51 action recognition, and ERA5 weather forecasting.
1 Introduction
Figure 1: Mamba-ND  outperforms Transformers while significantly reducing the number of parameters. On ImageNet-1k, we compare against ViT [12]. On HMDB-51[22], we compare against Video-Swin [28]. On ERA5, we compare against Cli-ViT [34].
The design of flexible and scalable neural network architectures is fundamental to the success of deep learning across diverse domains. Convolutional neural networks [23] excel at handling continuous data such as images, audio, and video. Recently, they have been surpassed by Transformers [45], which process continuous data as a discrete sequence of patches [12]. Despite their superior performance on many tasks such as Image Classification and Video Action Recognition, Transformer-based models struggle to scale to larger patch sequence as they scale quadratically with respect to sequence length. Most recently, a special kind of State Space Model (SSM) termed as Mamba [15], has demonstrated stronger performance than Transformers while maintaining a linear complexity. However, the impressive performance of Mamba was shown on 1D text sequences. This leaves open the question whether Mamba can be effectively extended to multi-dimensional data such as images, video, or scientific datasets, which is the focus of this work.
Unlike convolution or self-attention operations, which can be computed in parallel across the ND input data, Mamba requires a specific ordering of the data. Determining such an order is not an easy task. Building on past work in architecture design, we could consider many choices. One naive approach is to flatten the data in row-major order. Intuitively, this is non-optimal because in this setting information only flows in one direction, which is suboptimal for modeling complex relationships over multi-dimensional data that do not have any default ordering. Drawing inspiration from early works on Bi-directional LSTM, another alternative is to process the sequence in two directions at each layer and aggregate the results. We call this method Bi-SSM. While this design, in principle, allows information exchange between two arbitrary patches, two patches adjacent to each other spatially may have a huge distance between them on the computation graph. A natural extension of this method, which we call ND-SSM, is to process the input in $2D$ directions, where $D$ is the dimension of the data, and aggregate the results.
Rather than picking orders in sequence, another possibility is to borrow inspiration from multi-head self-attention in transformers, wherein we can split the channels into multiple heads and let each head process an SSM in a different direction. This design is similar to ND-SSM but differs in that it has less computational burden within each layer.
At a high level, there is also the question of block-level design, which specifies how Mamba layers are organized. For example, even using the vanilla Mamba layer as a black box, one can still apply the bi-directional or N-directional design at the block level by processing the same input through multiple layers with different directions.
In this work, we conducted an extensive study on these possible design choices. Surprisingly, we find that simply alternating between three fixed row-major orderings is one of the best-performing strategies on both 2D and 3D data. Armed with these findings, we propose Mamba-ND, a surprisingly simple yet effective design to extend Mamba to multi-dimensional data. Mamba-ND does not introduce any complicated changes to 1D-SSM layers. By stacking 1D-SSM layers as black boxes and alternating the sequence order between each layer, Mamba-ND was able to surpass Transformer-based models on various tasks, including image classification, action recognition, and weather forecasting, with a lower parameter count. It also maintains a linear complexity with respect to input sequence length. We show these results in Figure 1.
In summary, our main contributions are as follows.
  • We proposed Mamba-ND , which extends SSM to higher dimension through simply alternating sequence odering across layers.
  • Compared with Transformers, our model was able to achieve stronger performance at lower parameter count. Compared with ViT, our model achieves +1.5 increase in ImageNet-1K classification accuracy while reducing the number of parameters by 20.7%. On HMDB-51 video classification task, Mamba-ND  our model achieves +0.9 increase compared with 3D Swin-Transformer while reducing the number of parameters by 39.0%. On ERA5 5.625 degree 72 hour weather forecast tasks, Mamba-ND  achieves +0.7 gain in Anomaly Correlation Coefficient (ACC) while reducing the number of parameters by 44.5% when compared with Cli-ViT.
  • We conducted extensive ablation studies on various more complicated approaches that extend SSM to higher dimension. We find that complicated designs do not necessarily translate into stronger performance. These results should help future work on adapting 1D Sequence models to higher dimension.
2 Related Works
Image Modeling.
Convolutional Neural Networks (CNNs) [23, 17, 20, 39, 21, 38] have historically been the state-of-the-art approach for image recognition tasks. These networks learn to extract hierarchical features from images by using interleaved convolution and pooling layers. While they have demonstrated strong performance, the receptive field of convolution kernels only grows linearly with the model depth. This limitation restricts the ability of CNNs to effectively capture global context.
More recently, Vision Transformers (ViTs) [12] have emerged as a strong candidate for vision tasks. Drawing inspiration from the field of natural language processing, these models first divide an input image into a grid of discrete patches and then process them as a 1D sequence. While they provide a global receptive field through self-attention mechanisms at every layer, such advantages come with a cost. In particular, they exhibit quadratic computational and memory complexity with respect to input length, making them challenging to scale to large images.
To address this limitation, several works have proposed hybrid architectures that incorporate attention mechanisms into CNNs [47, 5, 24] or introduce hierarchical designs into transformers [26, 25, 8, 10, 11, 29, 40, 46, 52]. Other approaches, such as PixelRNN [43], attempt to apply recurrent models, which are inherently designed for linear complexity and global context, to images.
There has also been considerable efforts to apply state space models to 2D images, such as 2D-SSM[2] and S4ND [33]. Our work falls within the this category, wherein we specifically explore ways to adapt selective state space models to images.
Video Modeling.
Modeling videos has long presented a significant challenge. Earlier works [4, 41] primarily focused on extending 2D ConvNets into the 3D domain. More recently, 3D Transformers [1, 28, 32] have demonstrated superior performance. Some of these works add temporal attention on top of existing 2D Vision Transformers, while others attempt to apply self-attention in a pure 3D fashion.
The unique characteristics of the temporal dimension in video data have inspired various video-specific high-performing designs. For example, some works leverage extracted optical flow features [6, 14] or motion vectors [50, 51]. Others, like SlowFast [13], employ specific architectural designs to account for the fact that pixel values change less in the temporal dimension than in the spatial dimensions. In contrast to these approaches, our goal is to devise a generic framework for modeling multi-dimensional data. Consequently, Mamba-ND treats videos as simple 3D arrays of RGB pixels, making no additional assumptions about the temporal structure. This is crucial because in certain use cases such as weather forecasting, the third dimension of the data represents an additional spatial dimension instead of a temporal one.
S4ND [33] is a prior attempt to model videos using state space models (S4 [16]) and has achieved competitive performance. It directly extends the S4 formulation to higher dimensions and leverages the time invariance constraint to parallelize computations as the outer product of three 1D convolutions. In contrast to S4ND, we aim to extend selective state space models, i.e. Mamba, to higher dimensions, which do not adhere to linear time invariance (LTI). As a result, the convolutional trick is not applicable, necessitating alternative designs.
Climate Modeling.
Climate modeling presents a significant challenge in the 3D domain. We specifically focus on global weather forecasting. Typically, this task is evaluated using the ERA5 reanalysis dataset [18]. Early works [48, 49] employed CNN architectures. Pangu-Weather [3] was the first to introduce Transformers to this domain. It adapted a 3D-Swin Transformer to handle atmospheric measurements in 3D spaces. ClimaX [34] extended the Vision Transformer to weather data and proposed an architecture called Cli-ViT. Using this architecture, it established a strong foundational model, called ClimaX, capable of various tasks such as forecasting and downscaling through large-scale pretraining on simulated data.
Using state space models for climate modeling remains an unexplored area. We are the first to apply SSMs in the 3D weather forecasting task.
3 Background
3.1 State Space Models
Recent state space models (SSMs) [16, 33] have shown superior performance on long sequences. Formally, SSMs model the input data using the following ordinary differential equation (ODE):
$h^{\prime}(t)$ $=Ah(t)+Bx(t)$ (1)
$y(t)$ $=Ch(t)+Dx(t)$ (2)
Here, $x(t)\in\mathbb{R}$ is a continuous input signal in the time domain, and $y(t)\in\mathbb{R}$ is a continuous output signal in the time domain. In modern SSMs, this ODE is approximated through discretization. One common discretization method is the zero-order hold (ZOH) rule, which gives the following difference equation:
$\bar{A}$ $=\exp(\Delta A)$ (3)
$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$ (4)
$h_{t}$ $=\bar{A}h_{t-1}+\bar{B}x_{t}$ (5)
$y_{t}$ $=Ch_{t}$ (6)
Early works such as S4 [16] and S4ND [33] assume linear time invariance (LTI). This constraint makes it possible to solve the above difference equation using a global convolution kernel. Selective state space models, i.e., Mamba [15], introduce time-varying parameters that do not follow the LTI assumption. In particular, $\Delta$, $B$, and $C$ become functions of the input signal $x(t)$. Consequently, from Equations 3, 4, 5 and 6, $\bar{A}$ and $\bar{B}$ also become dependent on time. This makes it hard to parallelize the computation. A hardware-aware optimization that makes use of associative scan is employed to mitigate this issue.
3.2 Mamba Layers
Mamba [15] proposed an implementation of selective state space model (sSSM) layer (Figure 3 column 1). It consists of a 1D convolution, an SSM kernel, and a residual connection. In each layer, the input sequence is first processed by the convolution operation and then by the SSM kernel. The result is added back to the input through the residual connection. While the Conv1D operation can be easily extended to multiple dimensions with ConvND operations, it is non-trivial to convert the SSM to multiple dimensions. S4ND [33] resolved this by formulating higher-dimensional SSM as a time-invariant system of partial differential equations (PDEs), which can be computed using global convolution. However, since Mamba is not a time-invariant system, this approach is not applicable. In this work, we discuss various alternative ways to extend SSM to higher dimensions by flattening the input data into 1D sequences.
4 Methodology
Figure 2: Architecture of Mamba-ND: In this figure we visualize Mamba-3D. Given 3D input data, we patchify it into $L$ patches. During this process, we maintain the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1-D Mamba layers that process the sequence in alternating orderings. In 3D space, we use the order H+H-W+W-T+T-. In 2D space, the sequence would be H+H-W+W-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream tasks.
Mamba-ND explores several approaches that adapt Mamba to multidimensional data. The key element of these designs is to devise a combination of sequence orderings to flatten the multidimensional data into 1-D sequences. Intuitively, some level of bidirectional or multidirectional design is required to allow information exchange between two arbitrary data points in multidimensional space. This can be achieved at two levels: layer level and block level. As mentioned in Section 3.2, a Mamba layer consists of a 1D convolution, an SSM kernel, and a residual connection. One example of layer-level design is to pass the output of the convolution to two independent SSM kernels and sum up the results (Figure 3, col 2). Contrary to layer-level designs, block-level design keeps the internal design of a Mamba layer unchanged. Instead, it tries to achieve bidirectionality on the block level. For example, one block-level design is to alternate between each axis from one layer to another and apply a bidirectional design for each axis (Figure 6, row 2). One possible benefit of layer-level design is that it keeps the optimized fused kernel of a Mamba layer and the memory access pattern of the underlying array unchanged.
Figure 3: Variations of SSM Layer Design. Col 1 represents the standard 1D SSM layer. Col 2 represents Bi-SSM, which adds bidirectionality in a similar fashion as LSTM. Col 3 represents ND-SSM block, which extends Bi-SSM to more directions. Col 4 represents multi-head SSM block inspired by multi-head attention in Transformers.
Figure 4: Visualization of 2D scan orderings. We visualize the set of possible all scan ordering on 2D data. Arrow indicates the scan order.
Figure 5: Different ways of arranging Mamba layers. The first row visualizes alternating-directional design. The second row visualizes bidirectional design. The third row visualizes quad-directional design.
1
Figure 5: Different ways of arranging Mamba layers. The first row visualizes alternating-directional design. The second row visualizes bidirectional design. The third row visualizes quad-directional design.
Figure 6: Visualization of various Scan-Factorization policies. Col 1: No factorization, there is only 1 sequence. Col 2: Factorizing the 3D sequence into $D$ 2D sequences, where $D$ is the length of a single dimension. Col 3: Factorizing the 3D sequence into $D^{2}$ 1D sequences.
4.1 Scan Orderings
Consider the input N-dimensional data $X$ of shape $D_{1}\times D_{2}\times\ldots\times D_{N}$, where $D_{i}$ is the length of data along the $i$th dimension. Let $L=\prod_{i=1}^{N}D_{i}$ be the total sequence length. There are a total of $L!$ possible ways of flattening $X$ into a 1D sequence. However, we only consider a subset of these choices, which we call the scan orderings. Formally, a scan ordering is obtained by permuting the order of axes of the original data $X$, and flattening it into a 1D sequence either in the forward or reverse direction. Since there are $N!$ permutations of $N$ axes and 2 possible directions, there are a total of $2N!$ scan orderings.
We denote a particular scan ordering $s$ as $(k_{1},k_{2},\ldots,k_{N})\pm$, which represents the unique ordering obtained by first permuting the axis order from $1,2,\ldots,N$ to $k_{1},k_{2},\ldots,k_{N}$, and then flattening the sequence in row-major order. The symbols $+$ and -- indicate whether the order of the final 1D sequence is in forward or reversed direction.
We focus on 2D data of shape $H\times W$ and 3D data of $T\times H\times W$. The 2D data has 4 possible orderings: $(HW)+$, $(HW)-$, $(WH)+$, and $(WH)-$. The 3D data has 12 possible orderings; examples include $(HWT)+$ and $(WHT)-$. To provide a concrete example, the ordering $(WH)-$ refers to first permuting the 2D data into a 2D array of $W\times H$, then flattening it into a 1D sequence in row-major order, and finally reversing the order of this 1D sequence.
For simplicity, we use $H$ to represent $WH$ and $W$ to represent $HW$ for 2D data. Similarly, we use $H$ to represent $TWH$, $W$ to represent $THW$, and $T$ for $HWT$ for 3D data. In this notation, $(WH)-$ becomes $H-$ and $(HWT)+$ becomes $T+$. We also use $L$ to represent $THW$ or $HW$, as this is the naive way of flattening 3D and 2D data to a 1D sequence without changing the memory layout. Notably, the last dimension will be traversed continuously. We visualize 2D scan orderings in Figure 4.
4.2 Adapting the Mamba Layer
We explore three possible alterations to the standard Mamba layer design, which are illustrated in Figure 3.
Bi-SSM
layer passes the output of the convolution layer to two independent SSM kernels, one in the forward direction and another in the reversed direction. This design draws inspiration from bidirectional LSTM [7], in which two LSTM layers of opposite directions are stacked together in a similar fashion.
ND-SSM
layer extends Bi-SSM by incorporating additional SSMs to accommodate other possible orderings. In the 2D case, there are four orderings $W+$, $W-$, $H+$, $H-$.
Multi-head SSM
layer is a mimic of the multi-head design in the Transformer. It splits an input sequence of dimension $D$ into $H$ sequences of dimension $D/H$, where $H$ is the number of orderings. Each of the heads is then passed to separate SSM kernels in respective orderings. In the 2D case, the orderings are $W+$, $W-$, $H+$, $H-$.
4.3 Arranging Mamba Layers
In addition to making direct changes to the internal structure of Mamba, one can also change the way in which the layers are organized to achieve multi-directionality. We illustrate these variations in Figure 6.
Alternating-Directional: H+H-W+W-T+T-
keeps the sequential ordering of Mamba layers and changes the direction of SSM in each layer in an alternating fashion. The ordering is H+H-W+W-T+T-
Bi-Directional: [H+H-][W+W-][T+T-]
adopts a design on the block level. In each block, the input is passed to two Mamba layers at opposite directions. The ordering is [H+H-][W+W-][T+T-], where each [.][.][ . ] denotes a bidirectional block consisting of two layers. To avoid confusion, we will explicitly refer to this method as [H+H-][W+W-][T+T-]. The term Bi-Directional will mostly be used for the Bi-SSM layer mentioned in Section 4.2.
Quad-Directional: [H+H-W+W-][T+T-]
builds on top of the [H+H-][W+W-][T+T-] design. It further groups the H and W directions. This design is inspired by works in Video Recognition such as [42], which factorize 3D convolution into a 2D operation on the spatial dimensions and a 1D operation in the temporal domain.
There are more possible ways to organize multi-directional blocks, but they generally follow a similar design. Crucially, it is important to note that while each layer has a specific ordering, the SSM kernel operates on a single flattened input sequence. This means that all these layers have a global receptive field.
4.4 Scan Factorization
In order to mitigate the quadratic complexity of the Transformer, prior works [19] factorize full 3D attention into three 1D attentions along each axis. While SSMs already achieve linear complexity, the sequence length is still quadratic in the length of a single dimension. In this work, we also explore various ways of factorizing an SSM scan into multiple smaller scans. For an input array of dimensions $T\times H\times W$, the standard approach is to flatten it into a single sequence of length $THW$. Alternatively, we can factorize it into $T$ sequences with length $HW$, or $TH$ sequences with length $W$. We visualize these factorization techniques in Figure 6. Since SSM only retains one copy of a state per sequence in the linear scan process, increasing the number of sequences in parallel actually leads to an increase in memory consumption and training time because more hidden states need to be materialized in the GPU memory. However, we note that these sub-sequences need not be computed in parallel. Hence a better implementation in the future may reduce the memory cost of this design. Thus, we still investigate the performance of this design.
4.5 Final Design
After extensive experiments, our final design uses the standard 1D-SSM layer and an alternating-directional: (H+H-W+W-T+T-) at the block level. We find that this simple design surprisingly outperforms more complicated ones. We provide more details of our experiments in Section 5.
Our overall image is shown in Figure 2. Given multi-dimensional input data, we first patchify it into a 1D sequence. During this process, we keep track of the original 3D structure of the input. This sequence is then passed through $K$ Mamba-ND blocks, each of which consists of a chain of 1-D SSM layers that process the sequence in alternating orderings. In 2D space, we use the order H+H-W+W-. In 3D space, we use the order H+H-W+W-T+T-. Finally, the sequence is reshaped back to its original 3D structure and passed to task-specific heads for downstream tasks.
5 Experiments
Settings
We aim to evaluate the effectiveness of Mamba-ND  on various multi-dimensional data tasks. Specifically, we use ImageNet-1K [9] for image classification, HMDB-51 [22] for action recognition, and ERA5 5.625-degree weather forecasting [18] for 3D tasks. ImageNet-1K is a large-scale dataset containing 1.2 million images across 1000 classes, HMDB-51 is an action recognition dataset comprising 7,030 video clips categorized into 51 action classes, and ERA5 consists of 3D atmospheric weather measurements, such as temperature and wind speed, across 13 pressure levels. We use the standard train and validation split for ImageNet, split1 of HMDB-51, and data from the years 1979-2016 for ERA5 as the training set, data from 2017 as the validation set, and data from 2018 as the test set.
Metrics
We measure top-1 accuracy for image classification and action recognition tasks. For climate forecasting, we report both the Residual Mean Squared Error (RMSE) and the Anomaly Correlation Coefficient (ACC), which is calculated as:
$ACC=\frac{\sum(f-c)(a-c)}{\sqrt{\sum(f-c)^{2}\sum(a-c)^{2}}}$ (7)
Here, $f$ represents the forecast, $c$ is the climatology (long time average measurement), and $a$ is the prediction target.
5.1 Image Classification
Following the approach of previous studies [12, 33, 17], Mamba-ND  is trained on the ImageNet-1K dataset for 300 epochs using the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $31e-$ 3. The results are presented in Table 1. Our model demonstrates superior performance compared to transformer-based models when operating under similar conditions, and it achieves results on par with the state-of-the-art state-space model, S4ND [12]. We compare our results with Hyena [37], ViT [12], and S4ND [33]. Notably, Mamba-ND-S shows a remarkable improvement of $+1.5$ in accuracy when compared to ViT, while simultaneously reducing the parameter count by 20.7%. This performance gap is consistent with prior research on 1D sequences [15], where Mamba consistently outperforms transformers with fewer parameters.
Table 1: ImageNet 1K Classification Results. We report Top 1 Accuracy on the validation set. Mamba-ND-S shows a remarkable improvement of $+1.5$ in accuracy when compared to ViT-B while reducing the parameter count by 20.7%
Model Image Size Params Acc.$\uparrow$
ViT-B 384 86M 77.9
ViT-L 384 307M 76.5
S4ND-ViT-B 224 86M 80.4
Hynea-ViT-B 224 88M 78.5
Mamba-2D-T 224 24M 79.2
Mamba-2D-S 224 63M 79.4
Mamba-2D-B 224 92M 80.2
5.2 Video Action Recognition
Prior works [4, 28] demonstrate strong performance on video datasets by adapting ImageNet-pretrained vision models to 3D tasks. Following their strategies, we inflate Mamba-2D to Mamba-3D. Since we adopted an alternating design, each layer only sees a 1D input sequence. This means the sizes of weights in Mamba-2D and Mamba-3D layers are identical, and we can directly load the Mamba-2D checkpoint. The only weight that have different shapes are the patch embeddings. We adopted a temporal patch size of 2, so we duplicated the ImageNet weights along the time dimension and divide the value by 2. To address the change in ordering from H+H-W+W- to H+H-W+W-T+T-, we simply append new layers for T+T- every other four layers.
We sample 32 frames at a frame interval of 2 from each video clip, which amounts to around 2 seconds of video. We use the AdamW optimizer with $\beta=(0.9,0.999)$ and a learning rate of $31e-$ 3. The learning rate on the backbone is multiplied by a factor of 0.1. We train our model with a global batch size of 64 for 50 epochs. We select Inception-I3D [4] and ConvNeXt-I3D [27] as our convolutional baseline. For transformer-based models, we select Video Swin Transformer [28].
We report the results in Table 2. Our method achieves a $+0.9$ accuracy compared to Video Swin Transformer while using 39% fewer parameters. Compared to the state-of-the-art ConvNet ConvNeXt, we achieve a $+5.5$ accuracy.
[…]
Model Acc.$\uparrow$ Params Memory Samples/s
ConvNeXt-I3D* 58.1 29M - -
S4ND-ConvNeXt-3D* 62.1 29M - -
Inception-I3D 49.8 25M - -
[…]
Video-Swin-T 53.0 30M 35GB 22.6
Video-Swin-S 58.1 54M 73GB 39.2
Mamba-2D 51.2 24M 12GB 39.2
Mamba-3D 59.0 36M 17GB 19.6
5.3 Global Weather Forecast
For weather forecasting tasks, we use Cli-ViT [34, 35, 36] and Pangu-Weather [3] as our baselines. Because these models were originally trained on terabytes of high-resolution data, typically around 1 degree, for a long time, we cannot directly compare them due to computational constraints. Instead, we use the 5.625-degree version of the ERA data across 7 pressure levels. Our baselines Cli-ViT are based on Vision Transformer, and Pangu-Weather is based on Swin-Transformer. For Cli-ViT, we use the official implementation from ClimaX [34]. For Pangu-Weather, we re-implement it in PyTorch. We train all models from scratch for 50 epochs, using a learning rate of 0.0005. The learning rate follows a cosine decay schedule with a warm-up period of 5 epochs. For Mamba-3D, we use a patch size of $2\times 2\times 2$.
We show the results in Table 3. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
Table 3: ERA5 5.625 Degree Weather Forecast Results. We report the Anomaly Correlation Coefficient (ACC) and RMSE on geopotential at the 500 hPa pressure level. Compared to Cli-ViT, Mamba-3D achieved a $+0.7$ ACC while reducing the parameter count by 44.5%.
[…]
Cli-ViT ViT 108M 467 89.3
Pangu Swin 50M 462 89.0
Mamba-3D SSM 50M 433 90.1
5.4 Meta Architectures
We perform extensive ablation studies on various design choices mentioned in Section 4. We show that the alternating-directional design is the simplest and most effective one among a wide range of possible choices.
Layer Design
We use ImageNet-1K and HMDB-51 classification as our benchmarks. We perform an ablation study on Mamba-2D and Mamba-3D. We adopted various designs mentioned in Section 4.2. Results show that our final alternating-directional design achieves stronger performance than all proposed layer-wise changes. We show results in Table 4. Compared with the 1D-Mamba baseline, in which no special design is adopted after naively flattening the multi-dimensional sequence into a 1D sequence, we achieved a $+4.8$ accuracy on ImageNet and a $+26.9$ accuracy on HMDB-51.
Additionally, we observe that most of the proposed multi-directional designs are able to outperform the naive and bi-directional baselines by considerable margins, suggesting the importance of multi-directional design.
Table 4: Ablation Study on Layer Designs. We report top-1 accuracy on the ImageNet-1K validation set. The Alt-Directional design is the top-performing one.
[…]
Alt-Directional Block Level 79.4 59.0
Multi-Head-SSM Layer-Level 77.6 51.5
ND-SSM Layer-Level 77.2 46.7
1D-SSM - 76.4 34.9
Bi-SSM Layer-Level 74.6 32.1
Layer arrangement
We perform comprehensive ablation studies on various organizations of SSM layers mentioned in Section 4.3 at the block level. In addition to three choices discussed in Section 4.3, we also experiment with a hex-directional design [H+H-W+W-T+T-], in which inputs of a block are processed by multiple layers at different ordering in parallel, and the results are summed together in the end. We show the results in Table 5. Contrary to intuition and the layer-level results, adding multi-directional design at the block level degrades the performance. We hypothesize that this behavior is caused by the reduced number of model depths. Since we keep the number of layers fixed across all designs, incorporating some level of multi-directional parallel processing will inevitably reduce the depths of the computation graph. We provide more discussions in the discussion section.
Table 5: Ablation Study on Layer Arrangement. We report top-1 accuracy on the HMDB-51 dataset. H+H-W+W-T+T- is the top-performing design.
Ordering HMDB-51$\uparrow$
H+H-W+W-T+T- 59.0
[H+H-W+W-T+T-] 38.3
[H+H-][W+W-][T+T-] 49.8
[H+H-W+W-][T+T-] 47.4
5.5 Scan Factorization
While factorizing the sequence in the current implementation of Mamba leads to considerable memory and runtime overhead, these costs are external to the designs themselves and may be patched in the future. For example, a scan of N sequences of length $L$ can be considered as a long scan of one sequence of $NL$ with $\bar{A}$ from Equation 5 set to zero in places where the scan operation moves from one short sequence to another. Hence, we find it useful to study the effects of various factorization techniques despite their inferior runtime efficiency.
We show such results in Table 6. We also report the memory cost on a single Nvidia A100 GPU. While all factorizations lead to worse performance, we find that 2D+1D factorization outperforms the 2D+3D setup, suggesting there may be merits of having certain layers dedicated to processing temporal correspondence. Because the cost of 1D factorization is high, we choose not to explore it further at this time.
Table 6: Ablation Study on Various Factorization Policies + indicates layer ordering. For example, the notation 2D+3D represents a 2D factorization layer and a 3D layer with no factorization arranged sequentially. In all studies, the total number of layers is fixed. We also show the training memory cost and the number of sequences. $B$ is the batch size, and $D$ is the length of a single dimension.
Factorization HMDB-51$\uparrow$ Mem Num.
1D + 1D + 1D 44.5 80GB $O(BD^{2})$
2D + 1D 55.8 77GB $O(BD^{2})$
2D + 3D 51.9 18GB $O(BD)$
3D 59.0 17GB $O(B)$
5.6 Video Specific Ablations
When initializing new SSM layers in T+T- direction, we can adjust the scale of $\Delta$ to control the initial temporal receptive field. We show such results in Table 7. We find that 1.0 is the optimal value, which coincides with the prior conclusion from S4ND [33].
Table 7: Ablation Study on the Scaling Factor of $\Delta$ for HMDB51 Experiments. We report the performance on HMDB-51 under different initializations of temporal layers (T+ and T-). $\Delta=1.0$ is the optimal choice.
$\Delta$ HMDB-51$\uparrow$
0.1 58.0
0.2 57.8
1.0 59.0
5.0 55.3
6 Discussion
Figure 7: Effective Receptive Field of Various Designs. Darkness indicates the sensitivity of the central patch of the output to each pixel of the input image, normalized to the range of $(0,1)$. All images are 224x224. We use ImageNet-1K pretrained checkpoints for these visualizations.
6.1 Effective Receptive Field (ERF)
To validate the significance of multi-directional design, we visualize the effective receptive field (ERF) [30] of Mamba-ND  in Figure 7. ERF is computed by setting the gradient of the center patch to one and backpropagating through the network. In particular, we visualize the ERF of ImageNet-1K pretrained weights of Mamba-ND  against the Bi-directional and Uni-directional baseline (Table 4, row 5 and Table 4, row 4). Sensitivity is measured by the whiteness of a pixel in the visualization. The Uni-Directional model exhibits a sharp cutoff and is insensitive to all patches after it in the flattened sequence. This is clearly undesirable in data where causal relations do not exist. The Bi-Directional model exhibits a global receptive field, but it is heavily biased in the horizontal direction. This is also undesirable because images can have vertical structures. The Multi-Directional model exhibits a more uniform pattern in the sensitivity visualization, which explains their superior performance.
6.2 Depths versus Widths
In Section 5.4, we showed that a complicated multi-directional design may not necessarily lead to improved performance over the alternating-directional baseline. We hypothesize that this result is related to the effective depth of the model.
Each of the block-level designs can be represented by a directed acyclic graph (DAG) of layers, where the direction of edges implies the computation order. In this interpretation, designs other than H+H-W+W-T+T- are trading tree depths for widths. Existing literature [44] shows that depth is more important than width in deep neural networks. Our results seem to reaffirm this conclusion.
7 Conclusion
Transformers have been the go-to choice in image, language, and video tasks in recent years, just like ConvNets and RNNs in earlier years. Mamba [31] presents itself as a competitive challenger to Transformers in 1D sequence modeling. However, due to their intricate formulations discussed in Section 4, it is not straightforward whether they will be equally competitive in multi-dimensional data. In this work, we proposed Mamba-ND  which successfully extends the strong performance of Mamba to multi-dimensional space. It was able to outperform Transformers on a variety of tasks with significantly fewer parameters, while reducing the quadratic complexity to a linear one. Through extensive experiments with alternative designs, we demonstrate the importance of multi-directional design in dealing with multi-directional data and highlight its advantages over uni-directional and bi-directional baselines. We hope these results will facilitate future research in multi-dimensional modeling.
8 Acknowledgement
We are grateful to Microsoft Research for supporting this research through their Accelerate Foundation Models Research program.